How do you simplify these numbers? -2+5(4+3r)
Answer by ikleyn(52787)

.
-2 + 5*(4+3r) = -2 + 5*4 + 5*3r = -2 + 20 + 15r = (-2 + 20) + 15r = 18 + 15r.